OpenStudy (frankypoo):

A: Use implicit differentiation to find the derivative dy/dx. B; Find the slope of the curve at the given point. cos(y)=x;(0,PI/2)

OpenStudy (frankypoo):

Please help me out, thanks :)

OpenStudy (bahrom7893):

Cos(y)=x -Sin(y) (dy/dx) = 1 dy/dx = -1/Sin(y)

OpenStudy (bahrom7893):

But, Cos(y)=x, so y = ArcCos(x), so dy/dx = 1/Sin(ArcCos(x))

OpenStudy (anonymous):

dx/dy=-siny dy/dx=-1/siny

OpenStudy (bahrom7893):

That's for part A.

OpenStudy (anonymous):

m=-1/sinpi/2=-1

OpenStudy (bahrom7893):

Part B: x = 0, y = pi/2 dy/dx = -1/Sin(y) dy/dx = -1/Sin(pi/2) = -1/1 = -1

OpenStudy (frankypoo):

slope is -1?

OpenStudy (frankypoo):

@bahrom7893 ?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

yes franky

OpenStudy (frankypoo):

Alright, thanks.